# Computing with D-Algebraic Sequences

### v1

@@ 3.2 Some closure properties and examples
4 L.h.s. D-algebraic sequences
In this section, we focus on two important subclasses of D-algebraic sequences. We show that sequences from both classes are zeros of difference polynomials that are linear in their highest shift terms.
@@ 4.1 Holonomic sequences
In [TTW24], Worrell and the author proposed an algorithm to convert any holonomic equation into an l.h.s. algebraic difference equation. We revisit this result and complete its proof.
@@
When working with holonomic equations or sequences, we often use the word “degree” to refer to the holonomic degree.
@@
Every holonomic sequence of order $l$ and degree $d$ is a zero of an l.h.s. difference polynomial of order at most $l+d$.
@@
This is a linear system of $d$ equations in $(n,n^{2},\ldots,n^{d})^{T}$. We claim that the matrix of the system has full rank due to the shifts involved in the $\gamma_{j,k}$’s. Indeed, each $\gamma_{j,k}$ has the following form:
@@
Let $\lambda_{0},\lambda_{1},\ldots,\lambda_{d-1}\in\mathbb{K}(s(n),\ldots,s(n+l+d-1))$ such that $\sum_{j=0}^{d-1}\lambda_{j}\,\Gamma_{j}=0$. The latter is a homogeneous linear system of linearly independent equations in the $\lambda$’s and can only have the trivial solution. To see that, fix $k$ and $k^{\prime}$, $1\leq k\neq k^{\prime}\leq d$. Notice that for $10\leq j_{1}\neq j_{2}\leq d-$ 1, $\gamma_{j_{1},k}$ and $\gamma_{j_{2},k}$ can be seen as two independent variables because $S_{l+j_{1}}$ and $S_{l+j_{2}}$ are two linearly independent vectors over $\mathbb{K}(s(n),\ldots,s(n+l+d-1))$. Thus, while it might be possible to find $\lambda_{0},\lambda_{1},\ldots,\lambda_{d-1}\in\mathbb{K}(s(n),\ldots,s(n+l+d-1))$ so that
$\lambda_{0}\,\gamma_{0,k}+\lambda_{1}\,\gamma_{1,k}+\cdots+\lambda_{d-1}\,\gamma_{d-1,k}=0,$ (19)
it is impossible for the same $\lambda$’s to verify
$\lambda_{0}\,\gamma_{0,k^{\prime}}+\lambda_{1}\,\gamma_{1,k^{\prime}}+\cdots+\lambda_{d-1}\,\gamma_{d-1,k^{\prime}}=0,$ (20)
unless $\lambda_{0}=\lambda_{1}=\cdots=\lambda_{d-1}=0$.
Hence (16) is non-singular and for all positive integers $k\leq d$, $n^{k}\in\mathbb{K}(s(n),\ldots,s(n+d+l-1))$. Finally, we plug these rational expressions into $\sigma^{d}(p)$. Since all the $n^{k}$’s are free of $s(n+l+d)$, the resulting difference polynomial is l.h.s. of order at most $l+d$. ∎
Our implementation of the algorithm from the proof of Theorem 4.2 is now part of the DalgSeq subpackage of NLDE. The corresponding procedure is HoloToSimpleRatrec.
@@
A Somos-like sequence is an integral sequence defined with a rational recursion. Using Theorem 4.2, one can generate a Somos-like sequence as follows:
@@
    Then use the algorithm from the proof of Theorem 4.2 to convert (21) into an l.h.s. difference polynomial.
@@
Catalan numbers are defined by the formula $C(n)\coloneqq\frac{1}{n+1}\binom{2n}{n}$, and satisfy the equation $(n+2)s(n+1)-(4n+2)s(n)=0$, which is holonomic. Using HoloToSimpleRatrec, we convert that equation into the rational recursion
@@
A noteworthy observation emerges from the geometrical view of Theorem 4.2. For the present example, with $\mathbb{K}=\mathbb{R}$, let us consider the surface defined by
@@
Figure 1 illustrates this geometry from the first quadrant of the $xy$-plane.
Figure 1: A geometric view of Theorem 4.2 for Catalan numbers.
@@ 4.2 $C^{2}$-Finite sequences
A $C$-finite sequence solves a linear recurrence equation with constant coefficients. A $C^{2}$-finite sequence is a solution to a linear recurrence equation with $C$-finite term coefficients [JPNP23]. As in the differential case with $D^{2}$-finite functions, $C^{2}$-finite sequences are D-algebraic. An implementation for computing algebraic difference equations from $C^{2}$-finite sequences is provided in our package under the name CCfiniteToDalg. The approach is similar to that of $D^{2}$-finite functions as described in [TT23]. In this subsection, we rely on a result from [CDBMP23] to establish the following theorem.
@@
If $(s(n))_{n}$ is $C^{2}$-finite such that the leading $C$-finite term coefficient of its defining equation is non-zero for all $n\geq N\in\mathbb{N}$. Then $(s(n))_{n\geq N}$ satisfies an l.h.s. algebraic difference equation.

### v2

@@ 3.2 Some closure properties and examples
We now deduce a result for sequence acceleration that follows from Corollary˜3.3.
Corollary 3.5.
Let $(s(n))_{n\in\mathbb{N}}$ be a D-algebraic sequence of order $r>2$. Assume that $\Delta^{2}s(n)\neq 0$ for sufficiently large $n$. The Aitken transformation of $(s(n))_{n\in\mathbb{N}}$, say $(t(n))_{n\in\mathbb{N}}$, of general term $t(n)\coloneqq s(n)-\frac{\left(\Delta s(n)\right)^{2}}{\Delta^{2}s(n)}$ is D-algebraic of order $r$.
Example 3.6 (An example of Aitken acceleration [Gri16]).
In his example of Aitken acceleration, Greffins considered a logical extension of Fibonacci’s rabbit problem in which two neighboring rabbit populations are, in addition to growing, competing for units of grassland. The probabilistic process that results from this scenario potentially gives rise to two random walks involving the Fibonacci numbers. We focus on the second, whose expectation is given by
$\mathbb{E}_{n}=-2\sum_{k=2}^{n}\frac{1}{F_{k}F_{k+1}},\,n\geq 2,$
where $F_{n}$ is the $n$th Fibonacci number. A concise description of the problem was given in terms of an equivalent process involving black and white discs and an urn (see [Gri16, second and third paragraphs]). The sequence of interest here is $S(n)=-\frac{1}{2}\mathbb{E}_{n}$, which is convergent as a consequence of the exponential-type behavior of the Fibonacci numbers. The Fibonacci sequence $(F_{n})_{n\in\mathbb{N}}$ is a D-algebraic zero of the difference polynomial $p\coloneqq\sigma^{2}(x)-\sigma(x)-x$, with initial values $F_{0}=0,\,F_{1}=1$. We use closure properties to find an algebraic difference equation satisfied by $(S(n))_{n\in\mathbb{N}}$. From that equation, we compute an ADE satisfied by the Aitken transformation of $(S(n))_{n\in\mathbb{N}},$ which we denote $(T(n))_{n\in\mathbb{N}}$.
  1. 1.
    $(u(n))_{n\in\mathbb{N}}\coloneqq\left(\frac{1}{F_{n}\,F_{n+1}}\right)_{n\geq 2}$ satisfies the following second-order quadratic algebraic difference equation.
    $s\!\left(n\right)s\!\left(n+1\right)+2s\!\left(n\right)s\!\left(n+2\right)+s\!\left(n+1\right)^{2}-s\!\left(n+1\right)s\!\left(n+2\right)=0,$ (29)
    with initial terms $u(0)\coloneqq\frac{1}{2},\,u(1)\coloneqq\frac{1}{6}.$ Here, the corresponding dynamical system is similar to that of the reciprocal, with the difference that the function $f$ has the form $1/(w_{1}\,w_{2}),$ where $\sigma(w_{1})=w_{2}$.
  2. 2.
    $(S(n))_{n\in\mathbb{N}}$ satisfies a quadratic ADE of order $3$. We write the ADE as a rational recursion to highlight that the corresponding initial does not vanish at $(S(n))_{n\in\mathbb{N}}$ as it is not a $C$-finite sequence.
    $s\!\left(n+3\right)=-\frac{s\!\left(n\right)s\!\left(n+1\right)-3s\!\left(n\right)s\!\left(n+2\right)-2s\!\left(n+1\right)^{2}+6s\!\left(n+1\right)s\!\left(n+2\right)-2s\!\left(n+2\right)^{2}}{2s\!\left(n\right)-3s\!\left(n+1\right)+s\!\left(n+2\right)}$ (30)
    As initial terms we have $S(0)\coloneqq\frac{1}{2},\,S(1)\coloneqq\frac{2}{3}.$
  3. 3.
    $(T(n))_{n\in\mathbb{N}}$ is a solution of the following third-order ADE of degree $4$.
    $s(n)^{2}\,s(n+1)^{2}+6\,s(n)^{2}\,s(n+1)\,s(n+2)-8\,s(n)^{2}\,s(n+1)\,s(n+3)+9\,s(n)^{2}\,s(n+2)^{2}\\ -24\,s(n)^{2}\,s(n+2)\,s(n+3)+16\,s(n)^{2}\,s(n+3)^{2}-12\,s(n)\,s(n+1)^{2}\,s(n+2)+10\,s(n)\,s(n+1)^{2}\,s(n+3)\\ -16\,s(n)\,s(n+1)\,s(n+2)^{2}+44\,s(n)\,s(n+1)\,s(n+2)\,s(n+3)-24\,s(n)\,s(n+1)\,s(n+3)^{2}\\ -4\,s(n)\,s(n+2)^{3}+10\,s(n)\,s(n+2)^{2}\,s(n+3)-8\,s(n)\,s(n+2)\,s(n+3)^{2}+4\,s(n+1)^{3}\,s(n+2)\\ -4\,s(n+1)^{3}\,s(n+3)+8\,s(n+1)^{2}\,s(n+2)^{2}-16\,s(n+1)^{2}\,s(n+2)\,s(n+3)\\ +9\,s(n+1)^{2}\,s(n+3)^{2}+4\,s(n+1)\,s(n+2)^{3}-12\,s(n+1)\,s(n+2)^{2}\,s(n+3)\\ +6\,s(n+1)\,s(n+2)\,s(n+3)^{2}+s(n+2)^{2}\,s(n+3)^{2}=0.$ (31)
    Despite its relatively big size, (31) enables the computation of terms of $(T(n))_{n\in\mathbb{N}}$ by finding rational roots of relatively simple quadratic polynomials. For the three needed initial terms of $(T(n))_{n\in\mathbb{N}}$, we need $5$ consecutive terms of $(S(n))_{n\in\mathbb{N}}$. Using $S(j),j=0,\ldots,4$ we get $T(0)\coloneqq\frac{7}{9},T(1)\coloneqq\frac{58}{75},T(2)\coloneqq\frac{743}{960}.$ Plugging these values into (31) for $s(n),s(n+1),$ and $s(n+2)$, respectively, yields the following factored quadratic equation to compute $T(3)$.
    $\left(108300X-83797\right)\left(10140X-7847\right)=0.$ (32)
    The desired next term is the root obtained with the positive square root of the discriminant; this is $T(3)\coloneqq\frac{7847}{10140}\approx 0.7738658777$, which is correct to $4$ decimal places with respect to the limit. This level of accuracy is achieved at the $9$th term of $(S(n))_{n\in\mathbb{N}}$.
    Table 1 in [Gri16] presents a comparative analysis of the convergence rates for $(S(n))_{n}$ and $(T(n))_{n}$, demonstrating an example of acceleration achievable through Aitken’s delta-squared process. Finding an ADE satisfied by the accelerating sequence allows for the computation of its terms independently, without needing to refer back to the original sequence. $\blacksquare$
4 D-algebraic rational recursions
In this section, we focus on two subclasses of D-algebraic sequences, namely holonomic and $C^{2}$-finite sequences. We show that generic sequences from both classes are D-algebraic zeros of rationalizing difference polynomials, defined as difference polynomials that are linear in their highest shift terms (see Definition˜2.3). As in the previous section, we concentrate on simple D-algebraicity, which corresponds to sequences entirely defined by an irreducible difference polynomial. By generic sequences in these subclasses, we want to exclude any behavior arising, for instance, from repeated values within the sequence terms, which may require polymorphic D-algebraicity where one of the irreducible difference polynomials is rationalizing. The power series coefficients of $\arctan$ (which exhibit repeated zeros) provide one such example, as does the D-algebraic continuation of Bernoulli numbers discussed in Example˜2.14.
Application-wise, we note that rational recursions establish a natural connection to automata construction with rational updates in theoretical computer science [ADD+13, BDSW17].
@@ 4.1 Holonomic sequences
In [TTW24], Worrell and the author proposed an algorithm to convert any holonomic equation into a rationalizing algebraic difference equation. We revisit this result and complete its proof.
@@
We remind that our focus is on D-algebraic zeros, in the sense that for any rationalizing difference polynomial $p$ obtained from a holonomic difference polynomial, our interest is on sequence solutions $(s(n))_{n}$ such that the set $\{n\in\mathbb{N},\,\mathcal{I}_{p}(s(n))=0\}$ is finite. For holonomic difference polynomials, this is clear, so all their solutions are D-algebraic in $\mathbb{D}_{s}(n).$ However, it requires some more work to construct a rationalizing difference polynomial in $\mathbb{D}_{s}$ from a holonomic sequence. We provide some ideas of sequences that we would like to avoid in the example below.
Example 4.1.
The sequence of power series coefficients of $\arctan(z)$, say $(a(n))_{n}$ satisfies the holonomic equation
$ns\!\left(n\right)+\left(n+2\right)s\!\left(n+2\right)=0,$ (33)
with initial terms $a(0)=0,a(1)=1$ (see [TTK22] and references therein for further details). Using the algorithm from [TTW24] to convert (33) into a rationalizing ADE (or rational recursion) yields
$s\!\left(n+3\right)=-\frac{s\!\left(n+1\right)\left(-s\!\left(n+2\right)+s\!\left(n\right)\right)}{s\!\left(n+2\right)+3s\!\left(n\right)}.$ (34)
However, $(a(n))_{n}$ is not a D-algebraic solution of this ADE since all its terms of even indices are zero, forcing its initial to be zero infinitely many times. Nevertheless, starting a sequence solution of (33) at $n\geq 1$ with two nonzero initial terms enables to define a holonomic solution of (33) that is a D-algebraic solution of (34).
On the other hand, it is relatively simple to deduce a difference polynomial satisfied by $(a(n))_{n}$ by considering its sequences of zero terms and nonzero terms separately. For the zero terms, given $n\in\mathbb{N}$, and $s(n)=0$, we want $s(n+2)=0$, so the equation is simply $s(n+2).$ For the nonzero terms, we convert the $2$-fold equation (33) in to a $1$-fold equation using the change of variables $n\rightarrow 2n+1,$ $s(2n+1)\rightarrow s(n)$ (see [TTK22]). We obtain $(2n+1)s(n)+(2n+3)s(n+1)=0$, which we convert into the rationalizing difference polynomial
$(s(n+1)+2s(n))s(n+2)+s(n+1)s(n).$ (35)
Now we interpret $s(n),s(n+1)$ and $s(n+2)$ as $s(n),s(n+2)$ and $s(n+4)$, respectively. This enables taking the index gap with the zero terms into account. Hence $(a(n))_{n}$ is a D-algebraic zero of
$s(n+2)\,\left((s(n+2)+2s(n))s(n+4)+s(n+2)s(n)+s(n+1)\right),$ (36)
which can also be written as $\sigma^{2}(x)\left((\sigma^{2}(x)+2x)\sigma^{4}(x)+\sigma^{2}(x)x+\sigma(x)\right)\in\mathcal{D}_{\sigma}(\mathbb{K},x)$. The term $\sigma(x)$ (or $s(n+1)$) is added to ensure that only one factor vanishes at every $n\in\mathbb{N}$. One can also obtain this difference polynomial using the “D-algebraic continuation method” exploited in Example˜2.14. $\blacksquare$
@@
When working with holonomic equations or sequences, we often use the word “degree” to refer to the holonomic degree. The order of the minimal holonomic difference polynomial satisfied by a sequence is the order of that holonomic sequence.
In Example˜4.1, the sequence of nonzero terms is a D-algebraic zero of a second-order rationalizing difference polynomial. This indicates that these terms can be entirely described by a rational recursion, despite the zero terms at odd indices. To avoid these case-by-case treatments of zeros of holonomic difference polynomials, we consider the concept of almost all for sequences defined by them. They have the particularity of using the full basis of solutions, unlike our previous example, where one term of the basis is killed by 0 in the initial terms of the sequence.
Definition 4.3.
Let $p\in\mathbb{D}_{s}(n)$ be a holonomic difference polynomial of order $l_{1},$ and $q\in\mathbb{D}_{s}$ of order $l_{2}$, $l_{2}\geq l_{1}$. Let $\tilde{p}=\sigma^{N+1}(p),$ where $N$ is the maximum nonnegative integer root of the polynomial coefficients in $p.$ Let $(\epsilon_{p}(n))_{n>N}$ be the solution of $\tilde{p}$ with symbolic initial terms
$\epsilon_{p}(j+N)=X_{j}\in\{X_{1},\ldots,X_{l_{1}}\},$
where the $X_{j}^{\prime}$s are viewed as distinct polynomial variables not belonging to $\mathbb{K}$. We say that almost all zeros of $p$ are D-algebraic zeros of $q$ if $(\epsilon_{p}(n))_{n>N}$ is a generic zero of $q$.
Note that in Definition˜4.3, $\epsilon_{p}(l_{1}+1),\ldots,\epsilon_{p}(l_{2}-1)$ are computed using $p$; and the generic zeros and D-algebraic zeros of $q$ coincide since the polynomials are free of $X_{j}$’s. In Example˜4.1, we have $N=0$, and the corresponding $(\epsilon(n))_{n>0}$ is given by
$\left(X_{1},X_{2},-\frac{X_{1}}{3},-\frac{X_{2}}{2},\frac{X_{1}}{5},\frac{X_{2}}{3},-\frac{X_{1}}{7},-\frac{X_{2}}{4},\frac{X_{1}}{9},\frac{X_{2}}{5},-\frac{X_{1}}{11},-\frac{X_{2}}{6},\frac{X_{1}}{13},\ldots\right).$ (37)
One verifies that this sequence can also be generated with (34) using the first three terms. The purpose of Definition˜4.3 is twofold: to highlight the dependence on the first terms and help demonstrate the non-triviality of the main result from [TTW24].
Definition 4.4.
We say that almost every holonomic sequence is a D-algebraic zero of a family of difference polynomials $\mathcal{F}$, if for any holonomic difference polynomial $p,$ almost all the zeros $p$ are D-algebraic zeros of some $q\in\mathcal{F}.$
Besides excluding particular solutions of holonomic equations, Definition˜4.3 and Definition˜4.4 also neglect irrelevant difference polynomials satisfied by holonomic sequences. For instance, if a holonomic sequence satisfies a difference polynomial $p\in\mathbb{D}_{s}$ of order $l$, then it also satisfies the rationalizing difference polynomial $q=\sigma^{l+2}(x)\sigma(p)+p$. However, the corresponding symbolic solution $(\epsilon_{p}(n))_{n>N}$ is not a generic zero of $q$, as it is a zero of $\mathcal{I}_{q}$.
@@
Almost every holonomic sequence of order $l$ and degree $d$ is a D-algebraic zero of a rationalizing difference polynomial of order at most $l+d$ in $\mathbb{D}_{s}$.
@@
This is a linear system of $d$ equations in $(n,n^{2},\ldots,n^{d})^{T}$. We claim that the matrix of the system has a generic full rank due to the shifts involved in the $\gamma_{j,k}$’s. Indeed, each $\gamma_{j,k}$ has the following form:
@@
Let $\lambda_{0},\lambda_{1},\ldots,\lambda_{d-1}\in\mathbb{K}(s(n),\ldots,s(n+l+d-1))$ such that $\sum_{j=0}^{d-1}\lambda_{j}\,\Gamma_{j}=0$. The latter is a homogeneous linear system of linearly independent equations in the $\lambda$’s and can only have the trivial solution.
To see that, notice that for $0\leq j_{1}\neq j_{2}\leq d-1$, $\gamma_{j_{1},k}$ and $\gamma_{j_{2},k}$ can be seen as two independent variables because $S_{l+j_{1}}$ and $S_{l+j_{2}}$ are two linearly independent vectors over $\mathbb{K}(s(n),\ldots,s(n+l+d-1))$.
Furthermore, one can always find indices $1\leq k_{1}\neq k_{2}\leq d$ such that the constant coefficients $c_{j,i,k_{1}}$ and $c_{j,i,k_{2}}$ (appearing in front of $s(n+j+i)$ for some $i$) are distinct. This distinction is vital; without it, the holonomic equation would simplify to a $C$-finite equation of the same order, contradicting our hypothesis. This specific characteristic implies that $\gamma_{j,k_{1}}$ and $\gamma_{j,k_{2}}$ are also effectively distinct variables, given that they generically represent two linearly independent homogeneous linear forms in $\mathbb{K}[s(n+j),\ldots,s(n+l+j)]$.
Thus, the homogeneous system in the $\lambda$’s can be represented by the following matrix:
$M_{\gamma}\coloneqq\begin{bmatrix}\gamma_{0,1}&\gamma_{0,2}&\ldots&\gamma_{0,d}\\ \gamma_{1,1}&\gamma_{1,2}&\ldots&\gamma_{1,d}\\ \vdots&\vdots&\ldots&\vdots\\ \gamma_{d-1,1}&\gamma_{d-1,2}&\ldots&\gamma_{d-1,d}\\ \end{bmatrix}$. (42)
This matrix $M_{\gamma}$ can be regarded as a square matrix where its $d^{2}$ components are distinct variable components, such as $(x_{1},x_{2},\ldots,x_{d^{2}})$. Alternatively, and perhaps more naturally, $M_{\gamma}$ can be viewed as a square matrix whose components are homogeneous linear forms with generic coefficients. In this perspective, the components within the same column share the same underlying set of variables, while those of two consecutive columns exhibit an overlapping structure.
Therefore, the determinant of $M_{\gamma}$ is nonzero, implying that $\lambda_{0}=\lambda_{1}=\cdots=\lambda_{d-1}=0$. It is essential to note that this analysis relies on purely symbolic computation involving shifts of $s(n)$. The value of $|M_{\gamma}|$ for a particular numerical value of $n$ is not the primary concern; rather, the focus is on establishing that $|M_{\gamma}|$ is not the zero-polynomial in $\mathbb{K}[s(n),\ldots,s(n+l+d-1)]$.
Hence (39) is non-singular, which implies that for all positive integers $k\leq d$, $n^{k}$ belongs to the field $\mathbb{K}(s(n),\ldots,s(n+d+l-1))$.
Now, let us consider $\sigma^{d}(p)$, viewed as a holonomic difference polynomial. We must explain why the $d$th shift of the leading polynomial coefficient of $p$ does not vanish when substituting the expressions of the $n^{k}$’s. Indeed, if it were to vanish, this would imply that the equations of the linear system solved to obtain the $n^{k}$ are linearly dependent with the polynomial coefficient of $s(n+d+l)$. This, however, is impossible because all coefficients of the linear system in (39) are homogeneous linear forms in $\mathbb{K}[s(n),\ldots,s(n+l+d-1)]$, each involving at least two ‘variables’. Such homogeneous linear forms cannot linearly combine to produce the constant coefficients that are present in the polynomial coefficient of $s(n+d+l)$ in $\sigma^{d}(p)$.
Furthermore, the symbolic zero $(\epsilon_{p}(j))_{j>N}$ cannot vanish the leading polynomial coefficient of $\sigma^{d}(p)$ since writing all its terms in terms of $X_{1},X_{2},\ldots,X_{l}$ and values of $n=j$, would simply evaluate that leading polynomial coefficient at $j$. And since $j>N,$ where $N$ is the maximum nonnegative integer roots of the polynomial coefficients of $p$, we can be sure that the leading polynomial coefficient is nonzero for all $j>N$.
Finally, since all the $n^{k}$’s are free of $s(n+l+d)$, the resulting difference polynomial is rationalizing with order at most $l+d$ in $\mathbb{D}_{s}$, and almost every zero of $p$ is one of its D-algebraic zeros. ∎
Our implementation of the algorithm from the proof of Theorem˜4.5 is now part of the DalgSeq subpackage of NLDE. The corresponding procedure is HoloToSimpleRatrec.
@@
A Somos-like sequence is an integral sequence defined with a rational recursion. Using Theorem˜4.5, one can generate a Somos-like sequence as follows:
@@
    Then use the algorithm from the proof of Theorem˜4.5 to convert (43) into a rationalizing difference polynomial.
@@
Catalan numbers are defined by the formula $C(n)\coloneqq\frac{1}{n+1}\binom{2n}{n}$, and satisfy the equation (n+2)s(n+1)(4n+2)s(n)=0(n+2)s(n+1)-(4n+2)s(n)=0, which is holonomic. Using HoloToSimpleRatrec, we convert that equation into the rational recursion
@@
A noteworthy observation emerges from the geometrical view of Theorem˜4.5. For the present example, with $\mathbb{K}=\mathbb{R}$, let us consider the surface defined by
@@
Figure˜1 illustrates this geometry from the first octant of the xyzxyz-space.
Figure 1: A geometric view of Theorem˜4.5 for Catalan numbers.
@@ 4.2 $C^{2}$-Finite sequences
A $C$-finite sequence solves a linear recurrence equation with constant coefficients. A $C^{2}$-finite sequence is a solution to a linear recurrence equation with $C$-finite term coefficients [JPNP23]. As in the differential case with $D^{2}$-finite functions, $C^{2}$-finite sequences are D-algebraic. A “slow” Groebner bases-based implementation for computing ADEs from $C^{2}$-finite sequences is provided in our package under the name CCfiniteToDalg. The approach is similar to that of $D^{2}$-finite functions as described in [TT23]. However, in this subsection, we establish a result similar to Theorem˜4.5 for this class of sequences. This result can also be seen as a consequence of the result from [CDBMP23].
As in the previous subsection, the concept of “almost every” can also be defined. However, for this part, we only concentrate on the algebraic manipulations. Before stating the main result, let us illustrate the idea on a generic first-order $C^{2}$-finite sequence with second-order $C$-finite term coefficients.
Let $(s(n))_{n}$ be a first-order $C^{2}$-finite sequence with $C$-finite coefficients $(c_{1}(n))_{n}$ and $(c_{0}(n))_{n}$, such that
$c_{1}(n)\,s(n+1)$ $+c_{0}(n)\,s(n)=0,$ (50)
$\alpha_{1,1}\,c_{1}(n+1)+\alpha_{1,0}\,c_{1}(n)=c_{1}(n+2),$ $\quad\,\alpha_{0,1}\,c_{0}(n+1)+\alpha_{0,0}\,c_{0}(n)=c_{0}(n+2)$ (51)
where $\alpha_{i,j}\in\mathbb{K},\,i,j\in\{0,1\}.$ In what follows, we assume that denominators in rational expressions are nonzero for some integers $n$ within the domain of interest.
Our idea is to eliminate the $c_{i}$’s from the shifts of (50) by incremental substitution. We start by eliminating $c_{1}(n)$ and its shifts. From (50) and its first shift, we have
$c_{1}\!\left(n\right)=-\frac{c_{0}\!\left(n\right)s\!\left(n\right)}{s\!\left(n+1\right)},\quad\,c_{1}\!\left(n+1\right)=-\frac{c_{0}\!\left(n+1\right)s\!\left(n+1\right)}{s\!\left(n+2\right)}.$ (52)
We now proceed with the elimination of $c_{0}(n)$. We consider the second shift of (50) with the corresponding substitutions using (51) and (52). We obtain
$c_{0}\!\left(n\right)=-\frac{s\!\left(n+1\right)\left(s\!\left(n+2\right)^{2}\alpha_{0,1}-s\!\left(n+1\right)s\!\left(n+3\right)\alpha_{1,1}\right)c_{0}\!\left(n+1\right)}{s\!\left(n+2\right)\left(s\!\left(n+1\right)s\!\left(n+2\right)\alpha_{0,0}-s\!\left(n\right)s\!\left(n+3\right)\alpha_{1,0}\right)}.$ (53)
Given that all shifts of $C$-finite coefficients can now be written as rational multiples of $c_{0}(n+1)$, it follows that their substitution in the third shift of (50) yields the product of $c_{0}(n+1)$ and a rational expression in the shifts of $s(n)$, in which $s(n+4)$ appears linearly. Hence, after canceling $c_{0}(n+1)$ we deduce the desired recursion:
$s\!\left(n+4\right)=\frac{\mathsf{N}}{\mathsf{D}}\coloneqq\frac{\mathsf{N}(s(n),\ldots,s(n+3))}{\mathsf{D}(s(n),\ldots,s(n+3))},$
where
$\mathsf{N}=-s\!\left(n+3\right)\left(-s\!\left(n+2\right)s\!\left(n\right)s\!\left(n+3\right)\alpha_{0,1}^{2}\alpha_{1,0}+s\!\left(n+1\right)^{2}s\!\left(n+3\right)\alpha_{0,0}\alpha_{0,1}\alpha_{1,1}+s\!\left(n+2\right)^{2}s\!\left(n+1\right)\alpha_{0,0}^{2}-s\!\left(n+2\right)s\!\left(n\right)s\!\left(n+3\right)\alpha_{0,0}\alpha_{1,0}\right),$ (54)
$\mathsf{D}=s\!\left(n+2\right)^{2}s\!\left(n\right)\alpha_{0,1}\alpha_{1,0}\alpha_{1,1}-s\!\left(n+2\right)s\!\left(n+1\right)^{2}\alpha_{0,0}\alpha_{1,1}^{2}-s\!\left(n+2\right)s\!\left(n+1\right)^{2}\alpha_{0,0}\alpha_{1,0}+s\!\left(n+1\right)s\!\left(n\right)s\!\left(n+3\right)\alpha_{1,0}^{2}.$ (55)
This suffices to prove our following proposition.
Proposition 4.8.
Almost every first-order $C^{2}$-finite sequence with second-order $C$-finite term coefficients is a D-algebraic zero of a rationalizing difference polynomial of order at most $5$.
Note that if some of the coefficients in the $C^{2}$-finite equation are nonzero constants, then in the third shift of that equation, not all $C$-finite coefficients (including constants) would be expressible as a rational function in the shifts of $s(n)$ multiplied by some $c_{i}(n+1),\,i\in\{0,1\}$. Therefore, one more shift would be required to find the rational recursion. This explains why the bound is $5$ and not $4$.
Example 4.9.
We take $(c_{1}(n))_{n}$ as a solution of the Fibonacci recurrence equation, and $(c_{0}(n))_{n}$ defined with the recursion $c_{0}(n+2)=2\,c_{0}(n+1)+3\,c_{0}(n)$. We have $\alpha_{1,1}=\alpha_{1,0}=1,$ $\alpha_{0,1}=2,$ and $\alpha_{0,0}=3$. Using (54) and (55), we deduce the rational recursion
$s\!\left(n+4\right)=-\frac{s\!\left(n+3\right)\left(6s\!\left(n+3\right)s\!\left(n+1\right)^{2}-7s\!\left(n+3\right)s\!\left(n\right)s\!\left(n+2\right)+9s\!\left(n+2\right)^{2}s\!\left(n+1\right)\right)}{s\!\left(n+3\right)s\!\left(n+1\right)s\!\left(n\right)+2s\!\left(n\right)s\!\left(n+2\right)^{2}-6s\!\left(n+1\right)^{2}s\!\left(n+2\right)}.$ (56)
The Gröbner bases method implemented in CCfiniteToDalg returns this same equation in about $85$ seconds in CPU time. $\blacksquare$
Let us now present the general algorithm. To simplify notations, we assume that the input data can be updated within the algorithm.
Algorithm 2 $C^{2}$-finite to D-algebraic rational recursion
  • $C^{2}$-finite equation $(p):c_{0}(n)s(n)+\cdots+c_{l}(n)s(n+l)=0,\,l>0$.
  • $C$-finite equations $(q_{j}):c_{j}(n+r_{j})=\alpha_{i,0}c_{j}(n)+\cdots+\alpha_{j,r_{j}-1}c(n+r_{j}-1),$ $j\in\{0,\ldots,l\}.$
A rationalizing difference polynomial of order at most $l+\sum_{j=0}^{l}r_{j}$ in $\mathbb{D}_{s}$.
  1. 1.
    Set $r_{l}\coloneqq r_{l}-1$, unless some $r_{j}=0$ and $c_{j}(n)\neq 0$.
  2. 2.
    For each $j$, $0\leq j\leq l$ do:
    • 2.1.
      For each $k$, $0\leq k\leq r_{j}-1,$ do:
      • 2.1.1.
        $c_{j}(n+k)\coloneqq\text{solve}((p),c_{j}(n+k)).$ //root of a univariate linear polynomial.
      • 2.1.2.
        $(p)\coloneqq\sigma((p))$;
      • 2.1.3.
        Update $(p)$ and the $(q_{i}),\,0\leq i\leq l,$ by substituting $c_{j}(n+k)$ as solved.
  3. 3.
    Let $r\coloneqq l+\sum_{j=0}^{l}r_{j}$.
  4. 4.
    Return $s(n+r)=\text{solve}((p),s(n+r))$” or its associated difference polynomial.
@@
Almost every $C^{2}$-finite sequence of order $l$ with $C$-finite term coefficients of order $r_{j},j=~0,\ldots,l$, is a D-algebraic zero of a rationalizing difference polynomial of order at most $l+\sum_{j=0}^{l}r_{j}.$